OpenStudy (anonymous):

I need help to solve: Q= p-q/2, for p This is what i got for it. = 2*Q= 2* p-q/2 = 2Q = 2/1 * p-q/2 = 2Q/-q = p-q/-q = 2Q/-q =p 0r p= 2Q/-q Is it correct, if not what did I do wrong?

OpenStudy (angela210793):

Q= p-q/2, for p multiply 2 both sides 2Q=p-q(add q both sides) 2Q+q=p

OpenStudy (anonymous):

p = Q + q/2...ur mistake was when u multiplied by 2. it becomes 2Q= 2p-q and not 2Q=2p - q/2

OpenStudy (anonymous):

thanks Kboy I got it now.

OpenStudy (anonymous):

thanks Angela for the help.

OpenStudy (anonymous):

and angela if im analysing ur name correctly, then its ur bday in 9days time..

OpenStudy (angela210793):

uw Tymaree :) @Kboy is it 12 today? if it is then yea my bday is in 9 days :):)

OpenStudy (anonymous):

yes it is in South Africa. njoy it

OpenStudy (angela210793):

Yea in Albania/Europe is too(is just tht i've lost the sense of time :P :P) Thnx :)